SIMULATION OF STOCHASTIC DIFFERENTIAL EQUATIONS

YOSHIHIRO SAITO^{1} AND TAKETOMO MITSUI^{2}

1Shotoku Gakuen Women’s Junior College, 1-38 Nakauzura, Gifu 500, Japan 2Graduate School of Human Informatics, Nagoya University, Nagoya 464-01, Japan

(Received December 25, 1991; revised May 13, 1992)

Abstract. A lot of discrete approximation schemes for stochastic differential equations with regard to mean-square sense were proposed. Numerical experiments for these schemes can be seen in some papers, but the efficiency of scheme with respect to its order has not been revealed. We will propose another type of error analysis. Also we will show results of simulation studies carried out for these schemes under our notion.

Key words and phrases: Numerical solution, stochastic differential equations, error analysis, order of convergence.